(-6x8y3)(7x3y)(-xy4)
Answer by lenny460(1073)

(-6x8y3)(7x3y)(-xy4)
I am gonna take your above expression as:
(-6x^8y^3)(7x^3y)(-xy^4)
Multiply:
-6x^8 * 7x^3 = -42x^11
y^3 * y = y^4

(-42x^11y^4)(-xy^4)
-42x^11 * -x = 42x^12
y^4 * y^4 = y^8

Answer:
42x^12y^8


obuong3@aol.com